OpenStudy (anonymous):

really stuck on this one :/ the seats at a local baseball stadium are arranged that teach row has 5 more seats than the row below it. if there are 4 seats in the first row, how many are in row 23?

OpenStudy (amistre64):

we have a difference of +5 between rows right?

OpenStudy (anonymous):

115

OpenStudy (amistre64):

r1 = 4 r2 = 4+5 r3 = 4+5+5 r4 = 4+5+5+5 = 4 + 5(3) rn = 4 + 5(n-1)

OpenStudy (anonymous):

4+5(x-1) should get you the answer so for row 23 we have 114

OpenStudy (anonymous):

115

OpenStudy (amistre64):

r23 = 4 +5(23-1) = 4 +5(22) = 4 +110 = 114

OpenStudy (amistre64):

its only 115 if you count them by hand and lose track :)

OpenStudy (anonymous):

Thank you so much!~~~!!